OpenStudy (anonymous):

simplify 6^1/3 times 36^1/3 A. 36 B. 6 C. square root (6) D. 3^ square root (6)

OpenStudy (anonymous):

6^1 equals 6 divided by 3 36^1 equals 36 divided by 3

OpenStudy (anonymous):

So you have 6/3 * 36/3 which can be simplified to 2*12

OpenStudy (anonymous):

I did something wrong

OpenStudy (anonymous):

Damnit

OpenStudy (anonymous):

Idk...sorry

OpenStudy (anonymous):

its okay :P

OpenStudy (math&ing001):

\[6^{1/3}*36^{1/3}=6^{1/3}*6^{2/3}=6^{\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }}\]

OpenStudy (math&ing001):

Can you finish ?

OpenStudy (math&ing001):

@rockstargirl17

OpenStudy (anonymous):

6^3/3 ?

OpenStudy (math&ing001):

Yes and 3/3=?

OpenStudy (anonymous):

it would equal to 1

OpenStudy (math&ing001):

Yeah ! So what would be the answer ?

OpenStudy (anonymous):

hmm it cant be 7 cause thats not the answer choices

OpenStudy (math&ing001):

Well it's not 7. ^ is the power sign. 6^1=6, 6^2=36, 6^3=6*6*6... I hope you get it

OpenStudy (anonymous):

oh okay so it would end up being 6

OpenStudy (anonymous):

A composition of reflections over parallel lines is the same as a __________. A. translation B. rotation C. glide reflection D. double rotation

OpenStudy (math&ing001):

Yes correct. I hope you understand the process though, cause it's more important than the answer. We used this properties: \[(x ^{a})^{b}=x ^{a*b}\] \[x ^{a}*x ^{b}=x ^{a+b}\]